OpenStudy (zzr0ck3r):

Let \(A\) be a countable orthonormal set of a Hilbert space \(H\). Prove the following. a) \(P_{\overline{\text{span A}}}(x) = \sum_{a\in A}\langle x,a\rangle a \ \text{ for all }x\in H\) b) \(\rho(x,\overline{\text{span A}}) = ||x||^2-\sum_{a\in A}|\langle x,a \rangle |^2 \ \ \ \ \forall x\in H\) c)If \(\alpha\) is a scalar-valued function on \(A\) s.t. \(\sum_{a\in A}|\alpha(a)|^2<\infty\), then the sum \(\sum_{a\in A}\alpha(a)a \) converges

OpenStudy (zzr0ck3r):

So there were things I needed to prove before I started this and I have done all of them, namely if \(A:=\{a_n|n\in\mathbb{N}\}\) is an orthonormal set in a Hilbert space \(H\), then t.f.a.e. (i) \(A\) is a orthonormal basis of \(H\). (ii) \(\overline{\text{span A}}=H\) (iii) For any \(x\in H\), one has \(x=\sum_{n=1}^\infty \langle x, a_n \rangle a_n\) (iv) for any \(x\in H\), one has \(||x||^2=\sum_{n=1}^\infty | \langle x, a_n \rangle |^2\) I also proved \(\langle x, a_n \rangle = 0 \ \forall n\in \mathbb{N} \implies \ x=0\)

OpenStudy (zzr0ck3r):

@eliassaab

OpenStudy (zzr0ck3r):

\(\rho(x,K)\) is the distance from x to set \(K\) \(P_k(x)\) is the projection of x onto \(K\)